OpenStudy (anonymous):

Jackson has a large collection of baseball and football trading cards. He has 124 baseball cards and 108 football cards. What is the relationship of the total number of cards to the number of baseball cards in Jackson's collection?

OpenStudy (solomonzelman):

total=124+108=132 let the number of baseball cards be x and number of cards in total be y. y=x-108 Not sure about how to say it, but total is 108 cards more than baseball

OpenStudy (anonymous):

I need the answer not the way to solve it

OpenStudy (solomonzelman):

Relationship as x-y graph, equation, or what?

OpenStudy (solomonzelman):

There are 108 more cards in total than in Baseball.

OpenStudy (anonymous):

Its about ratios... so i need the simplified ratio @SolomonZelman

OpenStudy (solomonzelman):

124/232 Answer: 31/58